Session: May 2000 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

\(Other\) (roma at Tue Jun 6 15:05:12 METDST 2000) Success
Every scan is recorded.

Effelsberg (A. Kraus at Wed Jun 7 10:02:03 METDST 2000) Minor failures
The scan from UT1105-1110 was lost due to computer problems.

Torun (K.Borkowski at Wed Jun 7 11:43:09 METDST 2000) Minor failures
Problems with cal signal for Tsys continued. PI should treat these measurements with caution.

Westerbork (Tony Foley at Wed Jun 7 13:49:59 METDST 2000) Minor failures
Started with 12 telescopes in array (RT5 out with unstable total power and RT8 out with helium leak.) After 13:12 RTC and RTD also switched out as the phase stability at 3.6cm over 3km was rotten! Delay steering problem from 19:21 to 19:50 RT7 was in the array, but had the wrong polarization at 3.6cm for the hybrid for the VLBI input.

Noto (C.Trigilio at Thu Jun 8 10:57:06 METDST 2000) Success
Source down in 4 scans (9:06, 10:06, 11:44, 12:16 UT).

Jodrell_Bank (Alastair Gunn at Fri Jun 9 11:07:13 METDST 2000) Success
No known problems.